Number of occurrence of a tossing four coins

1
Expert's answer
2022-06-12T15:02:50-0400

Every coin tossing has 2 possible outcomes: a Head (H) and aTail (T).


When tossing four coins, we have "2\cdot 2 \cdot 2\cdot 2=16" outcomes:

HHHH, HHHT, HHTH, HHTT,

HTHH, HTHT, HTTH, HTTT,

THHH, THHT, THTH, THTT,

TTHH, TTHT, TTTH, TTTT.


Answer: 16.